### v2

# Stability for Strichartz inequalities: Existence of minimizers

Abstract.
We study the quantitative stability associated with the adjoint Fourier restriction inequality, focusing on the paraboloid and two-dimensional sphere cases. We show that these Strichartz-stability inequalities admit minimizers attaining their sharp constants, provided that these sharp constants are strictly smaller than the corresponding spectral-gap constants. Furthermore, for the two-dimensional sphere case, we obtain the existence of minimizers.
Key words and phrases:
Fourier restriction, Strichartz inequality, quantitative stability, minimizer, concentration-compactness.
2020 Mathematics Subject Classification:
Primary 42B10; Secondary 49J40, 35B38, 35Q41.
This work was supported in part by the National Key R&D Program of China [Grant No. 2023YFC3007303], National Natural Science Foundation of China [Grant Nos. 12071052 & 12271501 & 12471232 & 12501130], and China Postdoctoral Science Foundation [Grant Nos. GZB20230812 & 2024M753436 & YJC20250892].
1. Introduction
The study of Fourier restriction phenomena plays a central role in harmonic analysis over the past decades. Roughly speaking, the well-known Paraboloid/Sphere Fourier Restriction Conjecture due to Stein asks the following question: for some given exponents $(p,q)$, do the associated $L^{p}\to L^{q}$ adjoint Fourier restriction inequalities hold? For the case $p=2$, this question was fully solved by Tomas-Stein [Tom75]. In this paper, we focus on these Tomas-Stein inequalities for the paraboloid and sphere, which are also known as Strichartz inequalities [Str77] since the space-time Fourier transform of PDE-solutions are supported on these surfaces.
For convenience, we introduce some necessary terminologies. Denoting the paraboloid/sphere surface $S\subset\mathbb{R}^{d+1}$ and the associated adjoint Fourier operator $E_{S}$, then the Tomas-Stein inequalities state that
(1) $\left\|E_{S}f\right\|_{L^{q}(\mathbb{R}^{d+1})}\leq C_{\sharp}\|f\|_{L^{2}(S)},\quad q:=2+4/d,\quad C_{\sharp}:=\sup_{0\neq f\in L^{2}(S)}\frac{\left\|E_{S}f\right\|_{L^{q}(\mathbb{R}^{d+1})}}{\|f\|_{L^{2}(S)}}.$
Here the surfaces are equipped with their natural measure, and further details can be seen later. For these Strichartz inequalities, we introduce the associated Strichartz deficit functional
$\delta_{S}(f):=C_{\sharp}^{2}\|f\|_{L^{2}(S)}^{2}-\left\|E_{S}f\right\|_{L^{q}(\mathbb{R}^{d+1})};$
and we call a nonzero function $f_{0}\in L^{2}(S)$ a maximizer for $C_{\sharp}$ if it satisfies $\delta_{S}(f_{0})=0$; in addition, for a manifold $\mathcal{M}\subset L^{2}(S)$, we introduce the following manifold distance
$\mathrm{dist}(f,\mathcal{M}):=\inf_{g\in\mathcal{M}}\left\|f-g\right\|_{L^{2}(S)}.$
There are natural questions whether these Strichartz inequalities (1) admit maximizers, and moreover, whether these Strichartz inequalities are stable in the sense that this Strichartz deficit functional $\delta_{S}(f)$ can control this distance from the maximizer manifold $\mathrm{dist}(f,\mathcal{M})$. In fact, the maximizers are conjectured to be Gaussian functions for the paraboloid case, and conjectured to be constant functions for the sphere case; while these two conjectures are only confirmed for the paraboloid with $d=\{1,2\}$ by Foschi [Fos07] and Hundertmark-Zharnitsky [HZ06], then confirmed for the two-dimensional sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$ by Foschi [Fos15]. For more progress on this maximizer problems we refer to the survey papers [FOeS17, NOeST23, OeS24]. Next, for the quantitative stability of Strichartz inequalities [^1] Indeed, the precompactness of maximizing sequences can give a qualitative stability result, further details on this aspect are referred to [BOeSQ20, DY24, CS12, FLS16, Kun03, Ram12, Sha09, Sha16] and the references therein., Duyckaerts-Merle-Roudenko[DMR11] have shown the coercivity of a relevant quadratic form for the paraboloid with $d=\{1,2\}$; and recently, Gonçalves-Negro[GN22] have shown the following quantitative stability estimates for the paraboloid and sphere cases [^2] They have also studied the wave equation case, which essentially corresponds to the cone case. See also the papers [Gon19a, Gon19b, Neg23a, Neg23b] for similar quantitative stability on Fourier restriction type inequalities. In a related direction, see [CNOeS24] for weighted-perturbation stability of maximizers in spherical extension inequalities.: there exist $\eta\in(0,1)$ and $C_{\sharp\sharp}>0$ such that
(2) $\delta_{S}(f)\geq C_{\sharp\sharp}\mathrm{dist}(f,\mathcal{M})^{2},$
under the condition that $\mathrm{dist}(f,\mathcal{M})<\eta\|f\|_{L^{2}(S)}$; and this condition can be removed if the corresponding maximizer conjecture is confirmed. Similarly, we call a nonzero function $f_{0}\in L^{2}(S)$ a minimizer for $C_{\sharp\sharp}$ if it can make the inequality (2) an equality.
However, the results in [GN22] do not give too much information on the stability constants $C_{\sharp\sharp}$. On the one hand, what is the optimal value of this stability constant $C_{\sharp\sharp}$; on the other hand, whether this value $C_{\sharp\sharp}$ can be attained for a minimizer. As a comparison, we recall the corresponding research progress on the Sobolev inequality: the maximizers are exactly Aubin-Talenti manifold according to the celebrated works [Aub76, Lie83, Ros71, Tal76], then the Sobolev-stability questions are raised by Brézis-Lieb [BL85] and the corresponding Sobolev stability estimate is established by Bianchi-Egnell [BE91], recently the explicit stability constant is investigated by Dolbeault-Esteban-Figalli-Frank-Loss [DEF+25] whose result is dimensionally sharp, and the existence of minimizers for stability constant is obtained by König [K2̈5]. For further details on this aspect, we refer to the survey papers [DEF+24, Fra24] and the references therein, see also [Chr14, Chr17, CLT23, CLT24, DSW25, FMP08, Fus15, FZ22] for quantitative stability in different settings. Here we emphasize that in the Sobolev setting, by the work of [K2̈5, Section 1.1], there are two crucial phenomena (two-peak and spectral-gap) which both play a role and thus lead to two compactness thresholds.
The main purpose of this paper is, based on the work [GN22] and motivated by the work [K2̈5], to further investigate the quantitative Strichartz stability inequalities. In this paper, we have established the following consequences:
  • for the sharp Strichartz stability constants $C_{\sharp\sharp}$ in (2), we compute the corresponding spectral-gap constants which directly give some upper bounds for $C_{\sharp\sharp}$;
  • different from the Sobolev setting, for the existence of minimizers for $C_{\sharp\sharp}$, there is only one compactness threshold (the two-peak phenomenon is essentially vanished);
  • for the existence of minimizers for $C_{\sharp\sharp}$, we discover an interesting fact that the sphere and paraboloid cases exhibit quite different behaviors: we obtain the existence of minimizers for the two-dimensional sphere, but only conditional existence of minimizers for the paraboloid.
Our main results are summarized as follows. Let $C_{*}$ and $C_{**}$ denote the sharp stability constants in (2) for the paraboloid and the two-dimensional sphere situation respectively, see also the precise expressions in the following (5) and (7). Then we have
Theorem 1.1 (Sphere-stability minimizer).
There exists a minimizer for $C_{**}$.
Remark 1.2.
In fact, our result implies that the minimizing sequence for $C_{**}$ cannot approach the maximizer manifold. Therefore, one cannot investigate the sharp constant $C_{**}$ by investigating the local behavior near the maximizer manifold.
Theorem 1.3 (Paraboloid-stability minimizer).
There exists a minimizer for $C_{*}$, provided that
(3) $C_{*}<\widetilde{C}_{d},\quad\widetilde{C}_{d}:=\frac{d^{2}+d+2}{(d+2)^{3}}2^{\frac{2}{d+2}}\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}.$
Note that the case $d\geq 3$ relies on the aforementioned Gaussian maximizer conjecture, which will be precisely stated later as Conjecture 1.5.
Remark 1.4.
We indeed have shown that $C_{*}\leq\widetilde{C}_{d}$ must hold, and in fact, the equality $C_{*}=\widetilde{C}_{d}$ will imply the noncompactness [^3] Note that this noncompactness does not imply the nonexistence of minimizers. of minimizing sequences for $C_{*}$. We mention that, for this condition (3), similar strict inequalities hold for the quantitative stability of Sobolev inequality [K2̈3] and planar isoperimetric inequality [BCH17], while their methods do not seem to work in our setting.
In the following two subsections, we present some precise terminologies on the paraboloid case and two-dimensional sphere case respectively. Meanwhile, we explain the main difficulties of our problem and the main methods of our proof.
1.1. Paraboloid adjoint Fourier restriction
We recall the Fourier transform and Schrödinger operator
$\widehat{f}(\xi):=\int_{\mathbb{R}^{d}}e^{-2\pi ix\xi}f(x)\mathrm{d}x,\quad e^{it\Delta}f(x):=\int_{\mathbb{R}^{d}}e^{2\pi ix\xi-4\pi^{2}it|\xi|^{2}}\widehat{f}(\xi)\mathrm{d}\xi.$
Then the paraboloid situation Tomas-Stein/Strichartz inequality states that
(4) $\left\|e^{it\Delta}f\right\|_{L^{q}(\mathbb{R}^{d+1})}\leq\mathbf{S}_{d}\|f\|_{L^{2}(\mathbb{R}^{d})},\quad q:=2+4/d,\quad\mathbf{S}_{d}:=\sup_{0\neq f\in L^{2}}\frac{\left\|e^{it\Delta}f\right\|_{L^{q}(\mathbb{R}^{d+1})}}{\|f\|_{L^{2}(\mathbb{R}^{d})}}.$
We define the standard Gaussian function $G(x):=e^{-\pi|x|^{2}}$ and the Gaussian maximizer manifold
$\mathcal{G}:=\left\{g:g=e^{it_{0}\Delta}\left[\lambda^{d/2}e^{2\pi i(x-x_{0})\xi_{0}}e^{-\pi|\lambda(x-x_{0})|^{2}}\right],\;(\lambda_{0},\xi_{0},x_{0},t_{0})\in\mathbb{R}_{+}\times\mathbb{R}^{3}\right\}.$
Conjecture 1.5 (Gaussian maximizer).
The maximizers for paraboloid Strichartz inequality (4) are exactly Gaussian-type functions $g\in\mathcal{G}$.
As mentioned above, this conjecture has been confirmed for dimensions $d=\{1,2\}$ and remains open for higher dimensions $d\geq 3$. Thus there holds $\mathbf{S}_{1}=12^{-1/12}$ and $\mathbf{S}_{2}=2^{-1/2}$. We define the following paraboloid Strichartz deficit functional and the Gaussian distance functional
$\delta(f):=\mathbf{S}_{d}^{2}\|f\|_{L^{2}(\mathbb{R}^{d})}-\left\|e^{it\Delta}f\right\|_{L^{q}(\mathbb{R}^{d+1})}^{2},\quad\mathrm{dist}(f,\mathcal{G}):=\inf_{g\in\mathcal{G}}\left\|f-g\right\|_{L^{2}(\mathbb{R}^{d})}.$
The work of Gonçalves-Negro [GN22, Theorem 1.2] shows that this deficit functional can control the Gaussian distance functional: more precisely, if the Gaussian maximizer Conjecture 1.5 holds, then there exists a constant $C_{*}>0$ such that
(5) $\delta(f)\geq C_{*}\mathrm{dist}(f,\mathcal{G})^{2},\quad C_{*}:=\inf_{0\neq f\in L^{2}}\frac{\delta(f)}{\mathrm{dist}(f,\mathcal{G})^{2}}.$
To establish the existence of minimizers for this paraboloid Strichartz stability inequality (5), as stated in [K2̈5], we have the following two natural obstacles
  • Spectral-gap: the minimizing sequences might converge to the maximizer manifold.
  • Two-peak: the minimizing sequences might consist of two maximizer profiles which are non-interacting in the limit sense.
Next we explain these two cases one by one.
First for the spectral-gap phenomenon, we recall some facts. To establish this quantitative stability result (5), Gonçalves-Negro [GN22, Section 2] has made use of the corresponding tangent space and its orthogonal complement space. Let $T_{G}\mathcal{G}$ be the tangent space of the manifold $\mathcal{G}$ at the Gaussian function $G$ and denote its orthogonal complement space as $T_{G}\mathcal{G}^{\perp}$. Then by direct calculation, see also [GN22, Equation (4-1)], there holds
$T_{G}\mathcal{G}=\mathrm{span}_{\mathbb{C}}\left\{e^{-\pi|x|^{2}},x_{1}e^{-\pi|x|^{2}},x_{2}e^{-\pi|x|^{2}},\cdots,x_{d}e^{-\pi|x|^{2}},|x|^{2}e^{-\pi|x|^{2}}\right\}.$
Here we define the following spectral-gap constant
$C_{SG}:=\inf_{h\in T_{G}\mathcal{G}^{\perp}}\frac{\delta^{\prime\prime}(G)[h,h]}{2\|h\|_{2}^{2}},\quad\delta^{\prime\prime}(f)[h,h]:=\frac{\partial^{2}}{\partial\varepsilon^{2}}\Big|_{\varepsilon=0}\delta(f+\varepsilon h).$
By this definition, if $h$ is a minimizer for $C_{SG}$ then one can conclude [^4] Notice that we times one-half in the definition of spectral-gap constant to make sure this fact holds true. In addition, in the Proof of Proposition 1.6, Step 3, we can show that this spectral gap constant $C_{SG}$ can be attained by taking the second order radial Hermite-Gaussian function.
$h\perp G\quad\Rightarrow\quad\delta(G+\varepsilon h)=C_{SG}\|h\|_{L^{2}(\mathbb{R}^{d})}^{2}\varepsilon^{2}+o(\varepsilon^{2}),$
thus one can directly see that $C_{*}\leq C_{SG}$. Indeed, Gonçalves-Negro [GN22, Proof of Theorem 1.2] has shown that $C_{SG}>0$. In this paper, we can further investigate some monotonicity to show the following Proposition 1.6, which guarantees that the spectral-gap obstacle cannot appear due to the the condition (3).
Proposition 1.6 (Paraboloid spectral-gap constant).
The explicit value of the spectral-gap constant is
$C_{SG}=\widetilde{C}_{d}=\frac{d^{2}+d+2}{(d+2)^{3}}2^{\frac{2}{d+2}}\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}.$
Second for the two-peak phenomenon, we define the paraboloid two-peak function $f_{\lambda}$ and two-peak constant $C_{TP}$ as follows
$f_{\lambda}(x):=G(x)+G_{\lambda}(x),\quad G_{\lambda}(x):=\lambda^{d/2}G(\lambda x),\quad C_{TP}:=\lim_{\lambda\to 0}\frac{\delta(f_{\lambda})}{\mathrm{dist}(f_{\lambda},\mathcal{G})^{2}}.$
By these definitions one can directly see that $C_{*}\leq C_{TP}$. In fact, we can compute that
Proposition 1.7 (Paraboloid two-peak constant).
The explicit value of the two-peak constant is
$C_{TP}=\left(2^{\frac{2}{d+2}}-1\right)\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}.$
In summary, if there holds the strict inequality $C_{*}<\min\{C_{SG},C_{TP}\}$, then the spectral-gap and two-peak both cannot happen, and thus it is expected that the existence of minimizers can be established. In fact, by some direct computation, we can prove that the two-peak vanishes in our Strichartz setting, which comes from the following result.
Proposition 1.8 (Paraboloid two-peak vanishing).
The spectral-gap constant and the two-peak constant satisfy the following relation
$C_{SG}<C_{TP}.$
Remark 1.9.
This conclusion reveals that, unlike the Sobolev setting studied by König in [K2̈3, K2̈5] where both the two-peak and spectral-gap must be surmounted, the two-peak obstacle vanishes within our Strichartz setting, thereby reducing the problem solely to overcoming spectral-gap.
Proof of Proposition 1.8.
The proof follows from direct computation. To show the desired conclusion, by extracting the common factor, it is enough to show
$2^{\frac{2}{d+2}}\frac{d^{2}+d+2}{(d+2)^{3}}>2^{\frac{2}{d+2}}-1\quad\Longleftrightarrow\quad 1-2^{-\frac{2}{d+2}}>\frac{d^{2}+d+2}{(d+2)^{3}}.$
Applying the inequality $1-e^{-x}\geq x-x^{2}/2$ with $x:=2\ln 2/(d+2)$, we can obtain
$1-2^{-\frac{2}{d+2}}\geq\frac{2\ln 2}{d+2}-\frac{2(\ln 2)^{2}}{(d+2)^{2}}.$
Then we can further conclude that
$1-2^{-\frac{2}{d+2}}-\frac{d^{2}+d+2}{(d+2)^{3}}$ $\geq(d+2)^{-3}\left[2\ln 2(d+2)^{2}-2(\ln 2)^{2}(d+2)-d^{2}-d-2\right]$
$=(2\ln 2-1)d^{2}+\left[8\ln 2-2(\ln 2)^{2}-1\right]d+[8\ln 2-4(\ln 2)^{2}-2]>0.$
This implies the desired result and completes the proof. ∎
1.2. Two-dimensional sphere adjoint Fourier restriction
In order to be consistent with existing literature, for the spherical measure $\sigma$, we denote the spherical Fourier transform
$\widehat{f\sigma}(x):=\int_{\mathbb{S}^{d-1}}f(\theta)e^{-ix\theta}\mathrm{d}\sigma(\theta),\quad x\in\mathbb{R}^{d}.$
Then the sphere situation Tomas-Stein/Strichartz inequality states that
(6) $\left\|\widehat{f\sigma}\right\|_{L^{q}(\mathbb{R}^{d})}\leq\mathbf{M}_{d}\|f\|_{L^{2}(\mathbb{S}^{d-1})},\quad q:=2+4/(d-1),\quad\mathbf{M}:=\sup_{0\neq f\in L^{2}}\frac{\left\|\widehat{f\sigma}\right\|_{L^{q}(\mathbb{R}^{3})}}{\|f\|_{L^{2}(\mathbb{S}^{2})}}.$
We define the constant maximizer manifold
$\mathcal{C}:=\left\{g:g=\lambda e^{iy\theta},\;(\lambda,y)\in\mathbb{R}_{+}\times\mathbb{R}^{3}\right\}.$
Conjecture 1.10 (Constant maximizer).
The maximizers for sphere Strichartz inequality (6) are exactly constant-type functions $g\in\mathcal{C}$.
As mentioned above, this conjecture has been confirmed for dimensions $d=3$, and remains open for dimension $d=2$ and higher dimensions $d\geq 4$. Thus there holds $\mathbf{M}=2\pi$. We define the following deficit functional of sphere-Strichartz inequality and the constant distance functional
$\delta_{*}(f):=\mathbf{M}_{d}^{2}\|f\|_{L^{2}(\mathbb{S}^{d-1})}-\left\|\widehat{f\sigma}\right\|_{L^{q}(\mathbb{R}^{d})}^{2},\quad\mathrm{dist}(f,\mathcal{C}):=\inf_{g\in\mathcal{C}}\left\|f-g\right\|_{L^{2}(\mathbb{S}^{d-1})}.$
The work of Gonçalves-Negro [GN22, Theorem 1.3] shows that this deficit functional can control the constant distance functional: more precisely, for dimension $d=3$, then there exists a constant $C_{**}>0$ such that
(7) $\delta_{*}(f)\geq C_{**}\mathrm{dist}(f,\mathcal{C})^{2},\quad C_{**}:=\inf_{0\neq f\in L^{2}}\frac{\delta_{*}(f)}{\mathrm{dist}(f,\mathcal{C})^{2}}.$
Similar to the paraboloid situation, there are two natural obstacles named spectral-gap and two-peak. Next we explain these two cases one by one.
First for the spectral-gap phenomenon, we recall some facts. Let $T_{1}\mathcal{C}$ be the tangent space of the manifold $\mathcal{C}$ at the constant function $1$ and denote its orthogonal space as $T_{1}\mathcal{C}^{\perp}$. Then, by some direct calculation, see also [GN22, Page 1125], there holds
$T_{1}\mathcal{C}=\mathrm{span}_{\mathbb{R}}\left\{1,i,ix_{1},ix_{2},ix_{3}\right\}.$
Here we define the following spectral-gap constant
$C_{SG*}:=\inf_{h\in T_{1}\mathcal{C}^{\perp}}\frac{\delta_{*}^{{}^{\prime\prime}}(1)[h,h]}{2\|h\|_{2}^{2}},\quad\delta_{*}^{{}^{\prime\prime}}(f)[h,h]:=\frac{\partial^{2}}{\partial\varepsilon^{2}}\Big|_{\varepsilon=0}\delta_{*}(f+\varepsilon h).$
By this definition, if $h$ is a minimizer for $C_{SG*}$ then one can conclude [^5] Notice that we times one-half in the definition of spectral-gap constant to make sure this fact holds true. And in the Proof of Proposition 1.11, Step 2, we can show that the spectral gap constant $C_{SG*}$ can be attained by taking the second order spherical harmonic function.
$h\perp 1\quad\Rightarrow\quad\delta(1+\varepsilon h)=C_{SG*}\|h\|_{L^{2}(\mathbb{R}^{d})}^{2}\varepsilon^{2}+o(\varepsilon^{2}),$
thus one can see that $C_{**}\leq C_{SG*}$. Indeed, Gonçalves-Negro [GN22, Proof of Theorem 1.3] has shown that $C_{SG*}>0$. In this paper, by investigating some monotonicity in the work of Gonçalves-Negro and combining some spectral perturbative calculation due to König [K2̈3], we can show the following Proposition 1.11, which directly guarantees that the spectral-gap obstacle cannot appear.
Proposition 1.11 (Sphere spectral-gap constant).
The explicit value of the spectral-gap constant is
$C_{SG*}=\frac{8\pi^{2}}{5}.$
Furthermore, the stability-constant satisfies
$C_{**}<\frac{8\pi^{2}}{5}.$
Second for the two-peak phenomenon, we define the sphere two-peak function and sphere two-peak constant as follows
$f_{y}(\theta):=1+e^{i\theta y},\quad C_{TP*}:=\lim_{|y|\to\infty}\frac{\delta(f_{y})}{\mathrm{dist}(f_{y},\mathcal{C})^{2}}.$
This two-peak constant is well-defined due to the rotation symmetry, and by these definitions one can directly see that $C_{**}\leq C_{TP*}$. In fact, we can compute that
Proposition 1.12 (Sphere two-peak constant).
The explicit value of the sphere-two-peak constant is
$C_{TP*}=\left(2-\sqrt{2}\right)\mathbf{M}^{2}=(2-\sqrt{2})4\pi^{2}.$
Finally, it is obvious that the two-peak phenomenon vanishes since $C_{SG*}<C_{TP*}$, and moreover $C_{**}<\min\{C_{SG*},C_{TP*}\}$. Therefore the aforementioned two obstacles are eliminated, and we thus can obtain the desired unconditional existence of minimizers Theorem 1.1.
1.3. Outline of the paper
We show the paraboloid result Theorem 1.3 in Section 2, and show the two-sphere result Theorem 1.1 in Section 3. In each section, we proceed three subsections to investigate the spectral-gap, two-peak, and existence of minimizers respectively.
2. Conditional existence of minimizers: paraboloid
2.1. Spectral-gap
In this subsection, we compute the spectral-gap constant $C_{SG}$ in Proposition 1.6, which comes from some more precise estimate based on [GN22].
First, we recall that Gonçalves and Negro [GN22, Pages 1119-1123] have shown the following results by using spherical harmonic functions and Lens transform: the minima of
$f\mapsto\|f\|_{2}^{-2}\delta^{\prime\prime}(G)[f,f]$
over all $T_{G}\mathcal{G}^{\perp}$ is the same as the minima over $T_{G}\mathcal{G}^{\perp}\cap L_{rad}^{2}(\mathbb{R}^{d})$; moreover for any function $f\in L_{rad}^{2}(\mathbb{R}^{d})$ with the following Laguerre polynomials expression
(8) $f(x):=\sum_{m\geq 0}a(m)L_{m}^{d/2-1}(2\pi|x|^{2})e^{-\pi|x|^{2}},$
where $L_{m}^{d/2-1}$ is the standard Laguerre polynomial, there holds
(9) $\delta^{\prime\prime}(G)[f,f]=2^{\frac{2-d}{2+d}}q^{-\frac{d^{2}}{2d+4}}\sum_{m\geq 2}\left[1-c_{d}(m)\right]|a(m)|^{2}L_{m}^{d/2-1}(0),$
where the constant
$c_{d}(m):=\frac{q}{2}\sum_{j=0}^{m}\binom{m+d/2-1}{m-j}\binom{m}{j}(1-2/q)^{2m-2j}(2/q)^{2j}.$
Indeed, they have shown the stability of Strichartz inequality by showing that there exists a uniform constant $\varepsilon>0$ satisfies
$c_{d}(m)<1-\varepsilon,\quad\forall m\geq 2.$
Now, we are going to show that the function $c_{d}(m)$ decreases when $m\geq 2$, and then compute the value $c_{d}(2)$ to obtain the desired spectral-gap constant.
Proof of Proposition 1.6.
The proof proceeds in three steps. We first prove the monotonicity of the coefficients $c_{d}(m)$ in the initial two steps, and finally compute the desired spectral-gap constant.
Step 1: decreasing of $c_{d}(m)$ for $d=\{1,2\}$. For the case $d=2$, one can direct compute that
$c_{2}(m)=\frac{2\cdot\binom{2m}{m}}{4^{m}}=\frac{2\cdot(2m)!}{4^{m}(m!)^{2}},\quad\Rightarrow\quad\frac{c_{2}(m+1)}{c_{2}(m)}=\frac{2m+1}{2m+2}<1.$
For the case $d=1$, one can conclude that
$c_{1}(m)=3^{1-2m}S_{m},\quad S_{m}:=\sum_{j=0}^{m}\binom{m}{j}\binom{m-1/2}{m-j}4^{m-j}.$
Recall that
$\binom{m}{j}=\frac{\Gamma(m+1)}{\Gamma(j+1)\Gamma(m-j+1)}=\frac{m!}{j!(m-j)!}$
and
$\binom{m-1/2}{m-j}=\frac{\Gamma(m+1/2)}{\Gamma(j+1/2)\Gamma(m-j+1)}=\frac{\frac{(2m)!\sqrt{\pi}}{4^{m}m!}}{\frac{(2j)!\sqrt{\pi}}{4^{j}j!}(m-j)!}$
Then we compute that
$S_{m}=\sum_{j=0}^{m}\frac{(2m)!}{(2j)![(m-j)!]^{2}}.$
Next, by the trigonometric function integration formula
$\int_{0}^{\pi}\cos^{2k}\theta\mathrm{d}\theta=\pi\frac{(2k)!}{4^{k}(k!)^{2}},$
one can use binomial theorem to obtain the following integral representation
$S_{m}$ $=\sum_{k=0}^{m}\frac{(2m)!}{(2m-2k)!}\frac{1}{(k!)^{2}}$
$=\sum_{k=0}^{m}\binom{2m}{2k}\frac{4^{k}}{\pi}\int_{0}^{\pi}\cos^{2k}\theta\mathrm{d}\theta$
$=\frac{1}{2\pi}\int_{0}^{\pi}\left[(2\cos\theta+1)^{2m}+(2\cos\theta-1)^{2m}\right]\mathrm{d}\theta.$
This directly gives the integral representation of $c_{1}(m)$ as follows
$c_{1}(m)=\frac{1}{2\pi}\int_{0}^{\pi}\left[\left(\frac{2\cos\theta+1}{3}\right)^{2m}+\left(\frac{2\cos\theta-1}{3}\right)^{2m}\right]\mathrm{d}\theta.$
Notice that $|2\cos\theta\pm 1|\leq 3$. Therefore, we conclude that $c_{1}(m)\geq c_{1}(m+1)$ for all $m\geq 2$, and thus the case $d=1$ is proved.
Step 2: decreasing of $c_{d}(m)$ for $d\geq 3$. In the following we focus on the case $d\geq 3$. As shown in [GN22, Page 1123], by using the Jacobi polynomials
$P_{m}^{\alpha,\beta}(x):=2^{-m}\sum_{k=0}^{m}\binom{m+\alpha}{m-k}\binom{m+\beta}{k}(x-1)^{m-k}(x+1)^{k},\quad\alpha>-1,\beta>-1,$
one can obtain the following identity
$c_{d}(m)=(1+2/d)y^{m}P_{m}^{(0,d/2-1)}(x),\quad y:=\frac{d-2}{d+2},\;\;x:=\frac{y+y^{-1}}{2}=\frac{d^{2}+4}{d^{2}-4}.$
Thus we investigate the following ratio
$\frac{c_{d}(m+1)}{c_{d}(m)}=y\frac{P_{m+1}^{(0,d/2-1)}(x)}{P_{m}^{(0,d/2-1)}(x)}=yr_{m},\quad r_{m}:=\frac{P_{m+1}^{(0,d/2-1)}(x)}{P_{m}^{(0,d/2-1)}(x)}.$
Next we will prove the following result which direct implies the desired decreasing conclusion
(10) $r_{m}<y^{-1},\quad\forall m\geq 2.$
By the 3-term recurrence relation for the Jacobi polynomials [Sze75, Section 4.5] as follows
$2(n+1)(n+\alpha+\beta+1)(2n+\alpha+\beta)P_{n+1}^{(\alpha,\beta)}(x)$
$=(2n+\alpha+\beta+1)\left[(2n+\alpha+\beta+2)(2n+\alpha+\beta)x+\alpha^{2}-\beta^{2}\right]P_{n}^{(\alpha,\beta)}(x)$
$\quad-2(n+\alpha)(n+\beta)(2n+\alpha+beta+2)P_{n-1}^{(\alpha,\beta)}(x),$
we can take $\alpha=0$ and $\beta=d/2-1$ to conclude
$r_{m}$ $=\frac{(2m+d/2)(2m+d/2+1)}{2(m+1)(m+d/2)}x+\frac{-(2m+d/2)(d/2-1)^{2}}{2(m+1)(m+d/2)(2m+d/2-1)}$
$\quad-\frac{m(m+d/2-1)(2m+d/2+1)}{(m+1)(m+d/2)(2m+d/2-1)}\frac{1}{r_{m-1}}.$
Here we denote this relation as follows
$r_{m}=a_{m}x+b_{m}-\frac{c_{m}}{r_{m-1}},\quad a_{m}>0,\;b_{m}<0,\;c_{m}>0.$
Thus we investigate the function
$f_{m}(t):=a_{m}x+b_{m}-\frac{c_{m}}{t},\quad t>0;\quad\Rightarrow\quad f_{m}^{\prime}(t)=\frac{c_{n}}{t^{2}}>0.$
By this monotonicity, to prove the desired (10), it remains to show the following two facts
(11) $f_{m}(y^{-1})<y^{-1},\quad r_{m}<y^{-1}.$
First, recalling the definition of $x$ and $y$, we can compute that
$f_{m}(y^{-1})-y^{-1}=-\frac{2(d^{3}+4d^{2}m+16dm^{2}+16dm-4d-16m)}{(d-2)^{2}(d+2m)(m+1)(d+4m-2)}\frac{d-2}{d+2}<0,\quad\forall d\geq 3,\;\forall m\geq 2.$
Second, noticing that $r_{0}=y^{-1}-2/(d-2)<y^{-1}$ and then applying the induction argument together with the monotonicity of $f_{m}$, we can conclude that
$r_{m}=f_{m}(r_{m-1})\leq f_{m}(y^{-1})<y^{-1}.$
Hence we obtain the desired results (11) and thus complete the proof of our first step.
Step 3: the spectral-gap constant. We recall the radial Hermite–Gaussian functions
$\widetilde{L}_{m}(x):=L_{m}^{d/2-1}(2\pi|x|^{2})e^{-\pi|x|^{2}},\quad\Rightarrow\quad\langle\widetilde{L}_{m},\widetilde{L}_{n}\rangle=2^{-d/2}L_{m}^{d/2-1}(0)\delta_{mn},$
and they form an orthogonal basis of $L_{rad}^{2}(\mathbb{R}^{d})$. Here we have used the inner product notation $\langle f,g\rangle:=\int f\bar{g}\mathrm{d}x$ and the Kronecker delta notation $\delta_{mn}$. Therefore, using the function in (8) and the identity (9), we can see that
$\frac{\delta^{\prime\prime}(G)[f,f]}{\|f\|_{2}^{2}}=\frac{2^{\frac{2-d}{2+d}}q^{-\frac{d^{2}}{2d+4}}\sum_{m\geq 2}\left[1-c_{d}(m)\right]|a(m)|^{2}L_{m}^{d/2-1}(0)}{2^{-d/2}\sum_{m\geq 0}|a(m)|^{2}L_{m}^{d/2-1}(0)}.$
In the previous steps we have shown that $c_{d}(m)$ decreases with respect to $m\geq 2$. Hence the weighted inequality gives that
$\frac{\delta^{\prime\prime}(G)[f,f]}{\|f\|_{2}^{2}}\geq\frac{2^{\frac{2-d}{2+d}}q^{-\frac{d^{2}}{2d+4}}\left[1-c_{d}(2)\right]}{2^{-d/2}},$
where the equality holds if and only if $f(x)=k\widetilde{L}_{2}(x)$. Finally, a direct computation shows that
$c_{d}(x)=\frac{d^{3}+4d^{2}+10d+4}{(d+2)^{3}},\quad C_{SG}=\frac{d^{2}+d+2}{(d+2)^{3}}2^{\frac{2}{d+2}}\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}.$
This gives the desired explicit value and thus we complete the proof. ∎
2.2. Two-peak
In this subsection, we compute the two-peak constant $C_{TP}$ in Proposition 1.7, which essentially comes from the profile decomposition conclusion, and this classical decomposition can be seen in the works such as [BV07, CK07, DFY25, DY23, Ker01, MV98].
Lemma 2.1 (Gaussian distance representation).
The Gaussian distance functional satisfies
$\mathrm{dist}(f,\mathcal{G})^{2}=\|f\|_{2}^{2}-m(f),\quad m(f):=\sup_{g\in\mathcal{G}_{1}}\left(\Re\int f\bar{g}\mathrm{d}x\right)^{2},\quad\mathcal{G}_{1}:=\left\{g\in\mathcal{G}:\|g\|_{2}=1\right\}.$
And for each function $f\in L^{2}(\mathbb{R}^{d})$, its Gaussian distance $\mathrm{dist}(f,\mathcal{G})$ can be attained.
Proof of Lemma 2.1.
The first part of this lemma follows from a direct expansion, and the second part can be shown by applying an approximate identity argument [Gra14, Section 1.2.4]. For the sake of brevity, we omit the detailed proof here. Further details can be found in [K2̈5, Lemma 2.2] and [DEF+25, Lemma 3.3]. ∎
Proof of Proposition 1.7.
By the definition of two-peak constant, this proof relies on the behavior of each terms as $\lambda\to 0$. We investigate these terms one by one.
Step 1: we show that
(12) $\|f_{\lambda}\|_{2}^{2}=2^{1-d/2}+2\lambda^{d/2}-d\lambda^{d/2+2}+o(\lambda^{d/2+2})\quad\text{as}\;\;\lambda\to 0.$
Indeed, by direct calculation one can obtain
$\|f_{\lambda}\|_{2}^{2}$ $=\|G\|_{2}^{2}+\|G_{\lambda}\|_{2}^{2}+2\int_{\mathbb{R}^{d}}G(x)G_{\lambda}(x)\mathrm{d}x$
$=2^{1-d/2}+2\lambda^{d/2}(1+\lambda^{2})^{-d/2}$
$=2^{1-d/2}+2\lambda^{d/2}-d\lambda^{d/2+2}+o(\lambda^{d/2+2}).$
This gives the limit behavior of $\|f_{\lambda}\|_{2}^{2}$ and completes the first step.
Step 2: we show that
(13) $\|e^{it\Delta}f_{\lambda}\|_{q}^{q}=\frac{1}{2}q^{-d/2}+o(1)\quad\text{as}\;\;\lambda\to 0.$
First, one can compute that
$e^{it\Delta}G_{\lambda}(x)=\frac{\lambda^{d/2}}{(1+4\pi it\lambda^{2})^{d/2}}\exp\left(-\frac{\pi\lambda^{2}|x|^{2}}{1+4\pi it\lambda^{2}}\right),$
and
$\left|e^{it\Delta}G(x)\right|=\frac{1}{(1+16\pi^{2}t^{2})^{d/4}}\exp\left(-\frac{\pi|x|^{2}}{1+16\pi^{2}t^{2}}\right).$
Integrating with $x$ and then $t$, these further give
$\int_{\mathbb{R}^{d}}\left|e^{it\Delta}G\right|^{q}\mathrm{d}x=q^{-d/2}(1+16\pi^{2}t^{2})^{-1},$
and
$\int_{\mathbb{R}}q^{-d/2}(1+16\pi^{2}t^{2})^{-1}\mathrm{d}t=\frac{1}{4}q^{-d/2}=\frac{1}{4(2+4/d)^{d/2}}.$
By the scaling symmetry, we can see that
$\left\|e^{it\Delta}G_{\lambda}\right\|_{q}^{q}=\frac{1}{4(2+4/d)^{d/2}}.$
Next, to show the desired result (13), we estimate the following error term
$R(\lambda):=\|e^{it\Delta}f_{\lambda}\|_{q}^{q}-\|e^{it\Delta}G\|_{q}^{q}-\|e^{it\Delta}G_{\lambda}\|_{q}^{q},$
which satisfies
$R(\lambda)\lesssim\int|e^{it\Delta}G|^{q-1}|e^{it\Delta}G_{\lambda}|\mathrm{d}x\mathrm{d}t+\int|e^{it\Delta}G|^{q-1}|e^{it\Delta}G_{\lambda}|\mathrm{d}x\mathrm{d}t=:I_{1}(\lambda)+I_{2}(\lambda).$
We are going to show that
$\lim_{\lambda\to 0}I_{1}(\lambda)=0,\quad\lim_{\lambda\to 0}I_{2}(\lambda)=0.$
Indeed, by the stationary phase method, one can see the following dispersive estimate
$\|e^{it\Delta}f\|_{L_{x}^{\infty}}\lesssim|t|^{-d/2}\|f\|_{L_{x}^{1}},$
which, by the interpolation with $L^{2}$-conservation law, further implies that
$\|e^{it\Delta}G\|_{L_{x}^{q}}\lesssim\langle t\rangle^{-\frac{d}{d+2}},\quad\|e^{it\Delta}G_{\lambda}\|_{L_{x}^{q}}\lesssim\lambda^{\frac{d}{d+2}}\langle\lambda^{2}t\rangle^{-\frac{d}{d+2}},\quad\langle\cdot\rangle:=(1+|\cdot|^{2})^{1/2}.$
Hence, by Hölder’s inequality, we obtain
$I_{1}(\lambda)\leq\int\|e^{it\Delta}G\|_{L_{x}^{q}}^{q-1}\|e^{it\Delta}G_{\lambda}\|_{L_{x}^{q}}\mathrm{d}t\lesssim\lambda^{\frac{d}{d+2}}\int_{\mathbb{R}}\langle t\rangle^{-\frac{d+4}{d+2}}\langle\lambda^{2}t\rangle^{-\frac{d}{d+2}}\mathrm{d}t.$
Then we divide the time region into two parts: if $t\leq\lambda^{-2}$, then $\langle\lambda^{2}t\rangle\sim 1$ and
$\int_{|t|\leq\lambda^{2}}\langle t\rangle^{-\frac{d+4}{d+2}}\langle\lambda^{2}t\rangle^{-\frac{d}{d+2}}\mathrm{d}t\sim\int_{|t|\leq\lambda^{2}}\langle t\rangle^{-\frac{d+4}{d+2}}\mathrm{d}t\lesssim 1;$
if $t\geq\lambda^{-2}$, then $\langle\lambda^{2}t\rangle\sim\lambda^{2}|t|$ and
$\int_{|t|\geq\lambda^{-2}}\langle t\rangle^{-\frac{d+4}{d+2}}\langle\lambda^{2}t\rangle^{-\frac{d}{d+2}}\mathrm{d}t\lesssim\lambda^{-\frac{2d}{d+2}}\int_{|t|\geq\lambda^{-2}}|t|^{-2}\mathrm{d}t\lesssim\lambda^{\frac{4}{d+2}}.$
In summary, we conclude
$I_{1}(\lambda)\lesssim\lambda^{\frac{d}{d+2}};$
and similarly we can deduce the following decay estimate
$I_{2}(\lambda)\lesssim\lambda^{\frac{d}{d+2}}.$
These give the desired conclusion $R(\lambda)\to 0$ as $\lambda\to 0$, and thus we obtain the desired result (13).
Step 3: we show that
(14) $\mathrm{dist}(f_{\lambda},\mathcal{G})=2^{-d/2}+o(\lambda^{d/2})\quad\text{as}\;\;\lambda\to 0.$
By the Gaussian distance representation Lemma 2.1, we investigate the item $m(f_{\lambda})$. Since $f_{\lambda}$ is real-valued radial decreasing function, by writing $g\in\mathcal{G}_{1}$ as $g=g_{1}+ig_{2}$ and the inequality
$\frac{\left(\Re\int f_{\lambda}\bar{g}\mathrm{d}x\right)^{2}}{\int|g_{1}|^{2}+|g_{2}|^{2}\mathrm{d}x}\leq\frac{(\int f_{\lambda}g_{1}\mathrm{d}x)^{2}}{\|g_{1}\|_{2}^{2}},$
one can use rearrangement inequality [LL01, Theorem 3.4] to see that
$m(f_{\lambda})=\sup_{g\in\mathcal{G}_{1}}\left(\Re\int f_{\lambda}\bar{g}\mathrm{d}x\right)^{2}=\sup_{\mu>0}\left(\int f_{\lambda}G_{\mu}\mathrm{d}x\right)^{2}\Big/\|G\|_{2}^{2}=2^{-d/2}\sup_{\mu>0}\left(\int f_{\lambda}G_{\mu}\mathrm{d}x\right)^{2}.$
Denote $H_{\lambda}(\mu):=\int f_{\lambda}G_{\mu}\mathrm{d}x$. By direct computations, one can obtain
$H_{\lambda}(\mu)=\langle G,G_{\mu}\rangle+\langle G_{\lambda},G_{\mu}\rangle=\left(\frac{\mu}{1+\mu^{2}}\right)^{d/2}+\left(\frac{\lambda\mu}{\lambda^{2}+\mu^{2}}\right)^{d/2};$
then conclude the following inequality
$H_{\lambda}(\mu)\leq\|G+G_{\lambda}\|_{2}\|G_{\mu}\|_{2}\leq\left(\|G\|_{2}+\|G_{\lambda}\|_{2}\right)\|G\|_{2}=2^{1-d/2},$
where the equality holds if and only if $\lambda=\mu=1$. On the other hand, there holds the identity $H_{\lambda}(\mu)=H_{\lambda}(\mu^{-1}\lambda)$, which implies that we only need to consider the case $\mu\in[\sqrt{\lambda},\infty)$ as we consider the value of $m(f_{\lambda})$. Now for each fixed $\lambda\in(0,1/2)$, we investigate
$\sup_{\mu\in[\sqrt{\lambda},\infty)}H_{\lambda}(\mu)=\sup_{\mu\in[\sqrt{\lambda},\infty)}\left[\left(\frac{\mu}{1+\mu^{2}}\right)^{d/2}+\left(\frac{\lambda\mu}{\lambda^{2}+\mu^{2}}\right)^{d/2}\right]=:\sup_{\mu\in[\sqrt{\lambda},\infty)}\left[S(\mu)+T(\lambda,\mu)\right].$
For each $\lambda$, due to the limit $\lim_{\mu\to\infty}H_{\lambda}(\mu)=0$, we see that the supremum of $\mu$ can be attained, which we denote as $\mu(\lambda)$. We claim that
(15) $\mu(\lambda)=1+2^{d/2}\lambda^{d/2}+o(\lambda^{d/2}).$
Then using this claim, we conclude
$m(f)^{1/2}$ $=2^{d/4}\left[S(1)+S^{\prime}(1)(\mu(\lambda)-1)+o(\mu(\lambda)-1)+T(\lambda,1)+o(T(\lambda,1))\right]$
$=2^{d/4}\left[2^{-d/2}+\left(\frac{\lambda}{\lambda^{2}+1}\right)^{d/2}+o(\lambda^{d/2})\right]$
$=2^{-d/4}+2^{d/4}\lambda^{d/2}+o(\lambda^{d/2}).$
Hence by Lemma 2.1 and the conclusion of our first step, we obtain the desired conclusion
$\mathrm{dist}(f,\mathcal{G})^{2}=2^{1-d/2}+2\lambda^{d/2}-\left[2^{-d/4}+2^{d/4}\lambda^{d/2}+o(\lambda^{d/2})\right]^{2}+o(\lambda^{d/2})=2^{-d/2}+o(\lambda^{d/2}).$
It remains to prove the previous claim (15), which relies on some uniform analysis and implicit function theorem. By direct computation one can see
$S^{\prime}(\mu)=\frac{d}{2}\frac{1-\mu^{2}}{(1+\mu^{2})^{2}}\left(\frac{\mu}{1+\mu^{2}}\right)^{\frac{d}{2}-1},\quad S^{\prime\prime}(1)=-\frac{d}{2^{1+d/2}}<0,$
and
$H_{\lambda}^{{}^{\prime}}(\mu)=\frac{d}{2}\frac{1-\mu^{2}}{(1+\mu^{2})^{2}}\left(\frac{\mu}{1+\mu^{2}}\right)^{\frac{d}{2}-1}+\frac{d}{2}\frac{\lambda(\lambda^{2}-\mu^{2})}{(\lambda^{2}+\mu^{2})^{2}}\left(\frac{\lambda\mu}{\lambda^{2}+\mu^{2}}\right)^{\frac{d}{2}-1}.$
These imply that $H_{\lambda}(\mu)$ decreases on $\mu\in(1,\infty)$ and thus $\mu(\lambda)\in[\sqrt{\lambda},1]$. On the other hand, for $\mu\in[\sqrt{\lambda},\infty)$ we have the uniform convergence estimate
$T(\lambda,\mu)\leq\left(\frac{\lambda\mu}{\mu^{2}}\right)^{d/2}\leq\left(\frac{\lambda}{\mu}\right)^{d/2}\leq\lambda^{d/4}\to 0\quad\text{as}\;\;\lambda\to 0.$
We then show that for arbitrary $\varepsilon>0$ small enough, there exists $\eta$ such that if $\lambda\in(0,\eta)$ then $\mu(\lambda)\in(1-\varepsilon,1+\varepsilon)$. For arbitrary $\varepsilon>0$, since the function $S$ has unique max-value on $\mu=1$, we define
$\eta_{1}:=S(1)-\max\left\{\sup_{\mu\leq 1-\varepsilon}S(\mu),\sup_{\mu\geq 1+\varepsilon}S(\mu)\right\},\quad\eta:=\min\{\eta_{1},(\eta_{1}/4)^{4/d}\},$
then for all $\mu\notin(1-\varepsilon,1+\varepsilon)$, we have
$H_{\lambda}(\mu)\leq S(\mu)+\frac{\eta}{4}\leq S(1)-\frac{3\eta}{4}.$
The continuity of $S$ implies that there exists $\mu_{0}\in(1-\varepsilon,1)$ such that
$H_{\lambda}(\mu_{0})\geq S(1)-\frac{\eta}{4},$
which further implies $\mu(\lambda)\in(1-\varepsilon,1+\varepsilon)$. This shows that $\lim_{\lambda\to 0}\mu(\lambda)=1$. Next, we apply implicit function theorem to show the desired claim (15). Notice that
$F(\lambda,\mu):=H_{\lambda}^{{}^{\prime}}(\mu),\quad\Rightarrow\quad F(0,1)=S^{\prime}(1)=0,\;\;\frac{\partial}{\partial\mu}F(0,1)=S^{\prime\prime}(1)=-\frac{d}{2^{1+d/2}}\neq 0.$
Thus there exists unique function $\mu(\lambda)$ such that
$F(\lambda,\mu(\lambda))=0,\quad\mu(0)=1.$
When $\lambda\to 0$ and $\mu\sim 1$, we can use Taylor’s expansion to deduce
$T(\lambda,\mu)=\left(\frac{\lambda}{\mu}\right)^{\frac{d}{2}}\left[1+O(\lambda^{2}/\mu^{2})\right]^{-\frac{d}{2}}=\left(\frac{\lambda}{\mu}\right)^{\frac{d}{2}}+O(\lambda^{\frac{d}{2}+2}),$
and
$\frac{\partial}{\partial\mu}T(\lambda,\mu)=-\frac{d}{2}\lambda^{\frac{d}{2}}\mu^{-\frac{d}{2}-1}+O(\lambda^{\frac{d}{2}+2}),$
as well as
$S^{\prime}(\mu)=S^{\prime\prime}(\mu)(\mu-1)+O(|\mu-1|^{2})=-\frac{d}{2^{1+d/2}}(\mu-1)+O(|\mu-1|^{2}).$
Taking these expansion into the aforementioned equation $F(\lambda,\mu(\lambda))=0$, we obtain
$\frac{d}{2^{1+d/2}}[\mu(\lambda)-1]+O(|\mu(\lambda)-1|^{2})-\frac{d}{2}\lambda^{\frac{d}{2}}\mu(\lambda)^{-\frac{d}{2}-1}+O(\lambda^{\frac{d}{2}+2})=0,$
which directly gives the desired claim (15) by the fact that $\lim_{\lambda\to 0}\mu(\lambda)=1$.
Step 4: with the previous three steps finished, by taking the conclusions (12) and (13) and (14) into the following expression, we obtain
$\frac{\mathbf{S}_{d}^{2}\|f_{\lambda}\|_{2}^{2}-\|e^{it\Delta}f_{\lambda}\|_{q}^{2}}{\mathrm{dist}(f_{\lambda},\mathcal{G})^{2}}$ $=\frac{2^{-d/2}\left(2^{1-d/2}+2\lambda^{d/2}\right)-2^{-2/q}q^{-d/q}+o(1)}{2^{-d/2}+o(\lambda^{d/2})}$
$=\left(2^{\frac{2}{d+2}}-1\right)\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}+o(1).$
This implies the desired explicit value and thus completes the proof. ∎
2.3. Conditional compactness of minimizing sequences
In this subsection, based on the assumption (3) and spectral-gap Proposition 1.6 together with two-peak vanishing Proposition 1.8, we show the desired existence of minimizers for the paraboloid stability constant $C_{*}$.
Proof of Theorem 1.3.
The proof strategy is very similar to the proof of [K2̈5, Theorem 1.2]. The arguments are pretty long, but standard, as shown in [K2̈5, Section 4]. Here for simplicity, we briefly summarize the proof steps and omit the details.
  • For the minimizing sequence $f_{n}$, by the fact
    $\mathbf{S}_{d}^{2}-C_{SG}=2^{-\frac{d}{d+2}}\left(\frac{d}{d+2}\right)^{\frac{d^{2}}{2d+4}}\frac{d^{3}+4d^{2}+10d+4}{(d+2)^{3}}>0,$
    one can deduce $\|e^{it\Delta}f_{n}\|_{q}^{2}\geq c_{0}>0$ and then use the refined Strichartz estimate to obtain
    $f_{n}\rightharpoonup f_{0}\;\;\text{in}\;\;L^{2}(\mathbb{R}^{d}),\quad f_{0}\neq 0.$
    Then denoting $g_{n}:=f_{n}-f_{0}\rightharpoonup 0$, we obtain the profile decomposition conclusions
    $\|f_{n}\|_{2}^{2}=\|f_{0}\|_{2}^{2}+\|g_{n}\|_{2}^{2}+o(1),\quad\|e^{it\Delta}f_{n}\|_{q}^{q}=\|e^{it\Delta}f_{0}\|_{q}^{q}+\|e^{it\Delta}g_{n}\|_{q}^{q}+o(1).$
  • On the one hand, by investigating the infinity of parameters and using the fact that $m(f)$ can be attained for each $f$, we have
    $m(f_{n})=\max\{m(f_{0}),m(g_{n})\}+o(1).$
  • On the other hand, by using the minimizing sequence property (profile decomposition conclusions) and divide-two-part arguments, as well as scaling twice to reach mean-value and then investigating the monotonicity to deduce a contradiction to the sharp constant $C_{*}$, we can obtain
    $m(f_{0})=m(g_{n})+o(1).$
  • Therefore, by scaling and the divide-two-part arguments, as well as investigating monotonicity and the fact $C_{*}\leq C_{SG}<C_{TP}$ due to Proposition 1.8, which deduces a contradiction to the sharp constant $\mathbf{S}_{d}$, we conclude
    $g_{n}\to 0.$
  • Finally, by the strong convergence of $g_{n}$, it remains to show that the limit-denominator is non-zero which means $f_{0}\notin\mathcal{G}$. This indeed comes from the local asymptotic analysis [GN22, Theorem 2.1] and the assumption $C_{*}<C_{SG}$ due to Proposition 1.6, which means that the spectral-gap phenomenon cannot happen.
Here we have used the limit-version of [GN22, Theorem 2.1] which states that: if $\delta^{\prime}(G)=0$ and there exists $\rho>0$ such that
$\delta^{\prime\prime}(G)[h,h]\geq\rho\|h\|_{2}^{2},\quad\forall f\in T_{G}\mathcal{G}^{\perp},$
then for the sequence $f_{n}$ with $\mathrm{dist}(f_{n},\mathcal{G})\to 0$ one can obtain
$\mathrm{dist}(f_{n},\mathcal{G})^{-2}\delta(f)\geq\rho/2.$
Finally, this outline finishes the proof. ∎
3. Existence of minimizers: two-dimensional sphere
3.1. Spectral-gap
In this subsection, we compute the spectral-gap constant $C_{SG*}$ and its relation with the sphere Strichartz stability constant $C_{**}$ as stated in Proposition 1.11, which comes from some more precise estimate based on [GN22].
First, we recall that Gonçalves and Negro [GN22, Pages 1125-1126] have shown the following results by using spherical harmonic functions and Bessel functions: for any function $f\in L^{2}(\mathbb{S}^{2})$ with expression
(16) $f(\theta):=\sum_{k\geq 2}a(k)Y_{k}(\theta)\in(T_{1}\mathcal{C})^{\perp},$
where $Y_{k}$ is the standard real spherical harmonics with $\|Y_{k}\|_{L^{2}(\mathbb{S}^{2})}=1$, there holds
(17) $\delta_{*}^{{}^{\prime\prime}}(1)[f,f]=2\mathbf{M}^{2}\sum_{k\geq 2}|a_{k}|^{2}-c_{0}^{-1}\mathbf{M}^{2}\sum_{k\geq 2}c_{k}\left[4|a_{k}|^{2}+2(-1)^{k}\mathrm{Re}(a_{k})^{2}\right],$
where the constant
$c_{k}:=\int_{0}^{\infty}|J_{\frac{1}{2}}(r)|^{2}J_{\frac{1}{2}+k}(r)^{2}\mathrm{d}r,$
and $J_{k}$ is the Bessel function of the first kind. For further properties on Bessel functions and spherical harmonic functions, we refer to the chapters [Gra14, Appendix B] and [SW71, Chapter 4], see also the classical books [Wat44] and [M6̈6].
Proof of Proposition 1.11.
The proof proceeds in three steps. We prove the decreasing of the constant $c_{k}$ in the first step, and then we compute the spectral-gap constant in the second step, finally we show that the stability constant is strictly smaller than this spectral-gap constant in the third step.
Step 1: decreasing of $c_{k}$. We recall the spherical Bessel function $j_{k}(r)=\sqrt{\frac{\pi}{2r}}J_{k+\frac{1}{2}}(r)$ and the following relation
(18) $j_{k+1}(r)=\frac{2k+1}{r}j_{k}(r)-j_{k-1}(r),$
with
$j_{0}(r)=\frac{\sin r}{r},\quad j_{1}(r)=\frac{\sin r}{r^{2}}-\frac{\cos r}{r},\quad j_{2}(r)=\left(\frac{3}{r^{3}}-\frac{1}{r}\right)\sin r-\frac{3}{r^{2}}\cos r.$
Hence we obtain
$c_{k}=\frac{4}{\pi^{2}}\int_{0}^{\infty}\sin^{2}rj_{k}^{2}(r)\mathrm{d}r=\frac{2}{\pi^{2}}\int_{0}^{\infty}j_{k}^{2}(r)\mathrm{d}r+\frac{2}{\pi^{2}}\int_{0}^{\infty}\cos(2r)j_{k}^{2}(r)\mathrm{d}r=:A_{k}+B_{k}.$
For the item $A_{k}$, by the Weber–Schafheitlin integral formula [Wat44, Page 405] as follows
$\int_{0}^{\infty}t^{-1}J_{v}^{2}(at)\mathrm{d}t=\frac{1}{2v},\quad\mathrm{Re}v>0,$
we obtain
$A_{k}=\frac{1}{\pi}\int_{0}^{\infty}\frac{J_{k+1/2}^{2}(r)}{r}\mathrm{d}r=\frac{1}{(2k+1)\pi}.$
For the item $B_{k}$, we compute that
$B_{k}=\frac{\pi}{2}\int_{0}^{\infty}\frac{\cos(2r)}{r}J_{k+1/2}^{2}(r)\mathrm{d}r,$
and then we introduce the notation
$B_{k,\eta}:=\frac{\pi}{2}\int_{0}^{\infty}e^{-\eta r^{2}}\frac{\cos(2r)}{r}J_{k+1/2}^{2}(r)\mathrm{d}r,\quad\eta>0.$
By the asymptotic expansion of Bessel function [Wat44, Page 199], for $r\gg 1$, we obtain
$J_{k+1/2}(r)=\sqrt{\frac{2}{\pi r}}\cos\left(r-\frac{(k+1/2)\pi}{2}-\frac{\pi}{4}\right)+O(r^{-3/2}),\quad r\to\infty.$
Thus, using the dominated convergence theorem, we conclude
$\frac{\cos(2r)}{r}J_{k+1/2}^{2}(r)\in L^{1}(0,\infty),\quad\Rightarrow\quad B_{k}=\lim_{\eta\to 0}B_{k,\eta}.$
Then by the integral representation of Bessel function [Wat44, Page 48] as follows
$J_{k+1/2}(r)=\frac{(r/2)^{k+1/2}}{k!\sqrt{\pi}}\int_{-1}^{1}e^{irt}(1-t^{2})^{k}\mathrm{d}t,$
for each $\eta>0$, we can use the Fubini theorem to deduce
$B_{k,\eta}=\frac{1}{\pi(k!)^{2}2^{2k+1}}\int_{-1}^{1}\int_{-1}^{1}(1-t^{2})^{k}(1-s^{2})^{k}K_{k,\eta}(t+s)\mathrm{d}t\mathrm{d}s,$
where
$K_{k,\eta}(t):=\int_{0}^{\infty}e^{-\eta r}r^{2k}\cos(2r)e^{itr}\mathrm{d}r.$
By the Euler integral formula
$\int_{0}^{\infty}r^{k}e^{-\alpha r}\mathrm{d}r=\frac{k!}{\alpha^{k+1}},\quad\mathrm{Re}\alpha>0,$
we can write $\cos(2r)=(e^{2ix}+e^{-2ix})/2$ and then obtain
$K_{k,\eta}(t)=\frac{(2k)!}{2}\left[\frac{1}{(\eta-i(t+2))^{2k+1}}+\frac{1}{(\eta-i(t-2))^{2k+1}}\right],$
which satisfies
$\left|K_{k,\eta}(t)\right|\leq\frac{(2k)!}{2}\left(\frac{1}{|t+2|^{2k+1}}+\frac{1}{|2-t|^{2k+1}}\right).$
Hence we obtain the following estimate
$\left|(1-t^{2})^{k}(1-s^{2})^{k}K_{k,\eta}(t+s)\right|\lesssim(1-t^{2})^{k}(1-s^{2})^{k}\left(\frac{1}{|t+s+2|^{2k+1}}+\frac{1}{|2-t-s|^{2k+1}}\right)=:\widetilde{K}_{k}(t,s),$
and one can check that $\widetilde{K}_{k}(t,s)\in L^{1}([-1,1]^{2})$. Therefore, we can use the dominated convergence theorem to conclude
$B_{k}=\lim_{\eta\to 0}B_{k,\eta}=\frac{1}{\pi(k!)^{2}2^{2k+1}}\int_{-1}^{1}\int_{-1}^{1}(1-t^{2})^{k}(1-s^{2})^{k}K_{k}(t+s)\mathrm{d}t\mathrm{d}s,$
where
$K_{k}(t):=\lim_{\eta\to 0}K_{k,\eta}(t)=\frac{i(2k)!(-1)^{k}}{2}\left(\frac{1}{(t+2)^{2k+1}}-\frac{1}{(2-t)^{2k+1}}\right).$
Now, due to the fact $K_{k}(-t-s)=-K_{k}(t+s)$, we directly obtain
$B_{k}=0,\quad\Rightarrow\quad c_{k}=A_{k}+B_{k}=\frac{1}{(2k+1)\pi},$
which directly implies that the sequence $c_{k}$ is decreasing.
Step 2: the spectral-gap constant. By using the function in (16) and the identity (17), we can see that
$\frac{\delta_{*}^{{}^{\prime\prime}}(1)[f,f]}{\|f\|_{2}^{2}}\geq\frac{2\mathbf{M}^{2}c_{0}^{-1}\sum_{k\geq 2}(c_{0}-3c_{k})|a_{k}|^{2}}{\sum_{k\geq 0}|a(k)|^{2}},$
where the equality holds if and only if $a(k)=0$ for odd number $k$ and $a(k)$ is real-valued for even number $k$. In the previous step we have shown that $c_{k}$ decreases with respect to $k\geq 2$. Hence the weighted inequality gives that
$\frac{\delta_{*}^{{}^{\prime\prime}}(1)[f,f]}{\|f\|_{2}^{2}}\geq 2\mathbf{M}^{2}c_{0}^{-1}(c_{0}-3c_{2})=\frac{8\pi^{2}}{5},$
where the equality holds if we choose
$f(\theta)=aY_{2}^{0}(\theta)=a\sqrt{\frac{5}{16\pi}}(3\cos^{2}\theta-1),\quad a\in\mathbb{R}.$
In summary, we obtain the desired explicit value of sphere-spectral-gap constant $C_{**}=\frac{8\pi^{2}}{5}$.
Step 3: stability constant v.s. spectral-gap constant. By taking the function
$f_{\varepsilon}(\theta):=1+\varepsilon Y_{2}^{0}(\theta)=1+\varepsilon\sqrt{\frac{5}{16\pi}}(3\cos^{2}\theta-1),$
we investigate the corresponding Rayleigh quotient
$E(f_{\varepsilon}):=\frac{\delta_{*}(f_{\varepsilon})}{\mathrm{dist}(f_{\varepsilon},\mathcal{C})^{2}}.$
First for the item $\|f_{\varepsilon}\|_{L^{2}(\mathbb{S}^{2})}$, by orthogonality, we directly obtain
(19) $\|f_{\varepsilon}\|_{L^{2}(\mathbb{S}^{2})}^{2}=|\mathbb{S}^{2}|+\varepsilon^{2}.$
Second for the item $\mathrm{dist}(f,\mathcal{C})$, by direct computation, we investigate
$m(f_{\varepsilon})$ $=\frac{1}{4\pi}\sup_{x\in\mathbb{R}^{3}}\left[\int_{\mathbb{S}^{2}}(1+\varepsilon Y_{2}^{0}(\theta))e^{-ix\theta}\mathrm{d}\sigma(\theta)\right]^{2}$
$=4\pi\sup_{x\in\mathbb{R}^{3}}\left[\frac{\sin|x|}{|x|}-\varepsilon j_{2}(|x|)Y_{2}^{0}\left(\frac{x}{|x|}\right)\right]^{2}$
$=:4\pi\sup_{x\in\mathbb{R}^{3}}H_{\varepsilon}(x).$
We are going to show that for each $\varepsilon\in[0,\frac{2(1-\sin 1)\sqrt{5\pi}}{35})$, there holds
$\sup_{x\in\mathbb{R}^{3}}H_{\varepsilon}(x)=1,\quad\Rightarrow\quad m(f_{\varepsilon})=4\pi,$
which, by the constant distance representation Lemma 3.1, directly implies that
(20) $\mathrm{dist}(f,\mathcal{C})=\varepsilon^{2},\quad\text{for}\;\;\varepsilon\in\left[0,\frac{2(1-\sin 1)\sqrt{5\pi}}{35}\right).$
To show this fact, note that
$Y_{2}^{0}(\theta)\in\left[-\sqrt{5/(16\pi)},\sqrt{5/(4\pi)}\right],\quad\Rightarrow\quad|Y_{2}^{0}|\leq\sqrt{5/(4\pi)},\quad\forall\theta\in\mathbb{S}^{2}.$
Then writing spherical coordinate $x=r\theta$, for any $r>0$, we obtain
$|\sup_{\theta\in\mathbb{S}^{2}}H_{\varepsilon}(x)|\leq\left(\left|\frac{\sin r}{r}\right|+\varepsilon\sqrt{5/(4\pi)}|j_{2}(r)|\right)^{2}.$
Next, for this item, we divide the proof into two parts: the case $r\in[0,1]$ and the case $r\in[1,\infty)$.
Case 1: for $r\in[0,1]$, by the Taylor formula
$\sin r=r-\frac{r^{3}}{6}+\frac{r^{5}}{120}+R_{1},|R_{1}|\leq\frac{r^{7}}{5040};\quad\cos r=1-\frac{r^{2}}{2}+\frac{r^{4}}{24}+R_{2},|R_{2}|\leq\frac{r^{6}}{720},$
recalling the relation (18), we conclude
$|j_{2}(r)|$ $=\left|\frac{r^{2}}{15}-\frac{r^{4}}{120}+\left(\frac{3}{r^{3}}+\frac{1}{r}\right)R_{1}-\frac{3}{r^{3}}R^{2}\right|$
$\leq\frac{r^{2}}{15}+\frac{r^{4}}{120}+\left(\frac{3r^{4}}{5040}+\frac{r^{6}}{5040}\right)+\frac{3r^{4}}{720}$
$\leq\frac{r^{2}}{12}.$
And one can similarly obtain
$\left|\frac{\sin r}{r}\right|\leq 1-\frac{r^{2}}{7},\quad\forall r\in[0,1].$
Hence, for $0\leq\varepsilon<\frac{24\sqrt{5\pi}}{35}$, we conclude
$|\sup_{\theta\in\mathbb{S}^{2}}H_{\varepsilon}(x)|\leq\left[1-r^{2}\left(\frac{1}{7}-\frac{\varepsilon\sqrt{5/(4\pi)}r^{2}}{12}\right)\right]^{2}\leq 1,\quad\forall r\in[0,1],$
where the equality $|\sup_{\theta\in\mathbb{S}^{2}}H_{\varepsilon}(x)|=1$ holds if and only if $r=0$.
Case 2: for $r\in[1,+\infty)$, it is not hard to see the following estimates
$|\sin r/r|\leq 1/r\leq 1,\quad|\sin r/r|\leq\sin 1,$
and
$|j_{2}(r)|\leq\frac{3}{r^{3}}+\frac{3}{r^{2}}+\frac{1}{r}\leq 7.$
Hence, for $0\leq\varepsilon<\frac{2(1-\sin 1)\sqrt{5\pi}}{35}$, we conclude
$|\sup_{\theta\in\mathbb{S}^{2}}H_{\varepsilon}(x)|\leq\sin 1+7\varepsilon\sqrt{5/(4\pi)}<1,\quad\forall r\in[1,+\infty).$
In summary, due to the fact $\frac{2(1-\sin 1)\sqrt{5\pi}}{35}<\frac{24\sqrt{5\pi}}{35}$, for each $\varepsilon\in[0,\frac{2(1-\sin 1)\sqrt{5\pi}}{35})$ we see that
$\sup_{x\in\mathbb{R}^{3}}H_{\varepsilon}(x)=\lim_{|x|\to 0}H_{\varepsilon}(x)=1.$
Thus we obtain the desired conclusion (20). Third for the item $\|\widehat{f_{\varepsilon}\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4}$, we recall that
$\widehat{\sigma}(x)=4\pi\frac{\sin|x|}{|x|},\quad\widehat{Y_{2}^{0}\sigma}(x)=-4\pi j_{2}(|x|)Y_{2}^{0}(x/|x|),$
and
$\int_{\mathbb{S}^{2}}Y_{2}^{0}(\theta)\mathrm{d}\sigma(\theta)=0,\quad\int_{\mathbb{S}^{2}}\left(Y_{2}^{0}(\theta)\right)^{2}\mathrm{d}\sigma(\theta)=1,\quad\int_{\mathbb{S}^{2}}\left(Y_{2}^{0}(\theta)\right)^{3}\mathrm{d}\sigma(\theta)=\frac{\sqrt{5}}{7\sqrt{\pi}}.$
Thus we conclude
$\|\widehat{f_{\varepsilon}\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4}=I_{0}+I_{1}\varepsilon+I_{2}\varepsilon^{2}+I_{3}\varepsilon^{3}+I_{4}\varepsilon^{4},$
where
$I_{0}:=\|\widehat{\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4},\quad I_{1}:=4\int_{\mathbb{R}^{3}}|\widehat{\sigma}(x)|^{2}\mathrm{Re}\left(\widehat{\sigma}(x)\overline{\widehat{Y_{2}^{0}\sigma}}(x)\right)\mathrm{d}x,$
and
$I_{2}:=\int_{\mathbb{R}^{3}}\left[2|\widehat{\sigma}|^{2}|\widehat{Y_{2}^{0}\sigma}(x)|^{2}+4\left(\mathrm{Re}\left(\widehat{\sigma}(x)\overline{\widehat{Y_{2}^{0}\sigma}}(x)\right)\right)^{2}\right]\mathrm{d}x=6\int_{\mathbb{R}^{3}}(\widehat{\sigma})^{2}\left(\widehat{Y_{2}^{0}\sigma}(x)\right)^{2}\mathrm{d}x,$
as well as
$I_{3}:=4\int_{\mathbb{R}^{3}}|\widehat{Y_{2}^{0}\sigma}(x)|^{2}\mathrm{Re}\left(\widehat{\sigma}(x)\overline{\widehat{Y_{2}^{0}\sigma}}(x)\right)\mathrm{d}x=4\int_{\mathbb{R}^{3}}\left(\widehat{Y_{2}^{0}\sigma}(x)\right)^{3}\widehat{\sigma}(x)\mathrm{d}x,\quad I_{4}:=\|\widehat{Y_{2}^{0}\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4}.$
Writing spherical coordinate $x=r\theta\in[0,\infty)\times\mathbb{S}^{2}$, we can directly compute that
$I_{0}=4^{5}\pi^{5}\int_{0}^{\infty}\frac{\sin^{4}r}{r^{2}}\mathrm{d}r,\quad I_{1}=-4^{5}\pi^{4}\int_{0}^{\infty}\frac{\sin^{3}r}{r}j_{2}(r)\mathrm{d}r\int_{\mathbb{S}^{2}}Y_{2}^{0}(\theta)\mathrm{d}\sigma(\theta)=0,$
and
$I_{2}=6(4\pi)^{4}\int_{0}^{\infty}\sin^{2}rj_{2}^{2}(r)\mathrm{d}r,\quad I_{3}=-4^{5}\pi^{4}\int_{0}^{\infty}r\sin rj_{2}^{3}(r)\mathrm{d}r\int_{\mathbb{S}^{2}}\left(Y_{2}^{0}(\theta)\right)^{3}\mathrm{d}\sigma(\theta).$
Then, similar to Step 1, by some direct computation [^6] One can also use some software such as Mathematica to check these integrals., one can obtain
$\int_{0}^{\infty}\frac{\sin^{4}r}{r^{2}}\mathrm{d}r=\frac{\pi}{4},\quad\int_{0}^{\infty}\sin^{2}rj_{2}^{2}(r)\mathrm{d}r=\frac{\pi}{20},\quad\int_{0}^{\infty}r\sin rj_{2}^{3}(r)\mathrm{d}r=-\frac{\pi}{28},$
these integrals further give the following
$I_{0}=256\pi^{6},\quad I_{1}=0,\quad I_{2}=\frac{384\pi^{5}}{5},\quad I_{3}=\frac{256\sqrt{5}\pi^{9/2}}{49}.$
In summary, for small $\varepsilon\ll 1$, we have established the following three conclusions
  • $\|f_{\varepsilon}\|_{L^{2}(\mathbb{S}^{2})}^{2}=|\mathbb{S}^{2}|+\varepsilon^{2}$.
  • $\|\widehat{f_{\varepsilon}\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4}=256\pi^{6}+\frac{384\pi^{5}}{5}\varepsilon^{2}+\frac{256\sqrt{5}\pi^{9/2}}{49}\varepsilon^{3}+o(\varepsilon^{3})$.
  • $\mathrm{dist}(f_{\varepsilon},\mathcal{C})^{2}=\varepsilon^{2}$.
Therefore, as $\varepsilon\to 0$, we obtain that
$E(f_{\varepsilon})$ $=\frac{16\pi^{3}+4\pi^{2}\varepsilon^{2}-\left(256\pi^{6}+\frac{384\pi^{5}}{5}\varepsilon^{2}+\frac{256\sqrt{5}\pi^{9/2}}{49}\varepsilon^{3}+o(\varepsilon^{3})\right)^{1/2}}{\varepsilon^{2}}$
$=\frac{8\pi^{2}}{5}-\frac{8\sqrt{5}\pi^{3/2}}{49}\varepsilon+o(\varepsilon).$
This gives the desired result $C_{**}<C_{SG*}=8\pi^{2}/5$, and thus we have finished the proof. ∎
3.2. Two-peak
In this subsection, we compute the two-peak constant $C_{TP*}$ in Proposition 1.12, which essentially comes from the profile decomposition conclusion, and this classical decomposition can be seen in the works such as [CS12, FLS16, FS24].
Lemma 3.1 (Constant distance representation).
The constant distance functional satisfies
$\mathrm{dist}(f,\mathcal{C})^{2}=\|f\|_{2}^{2}-m(f),\quad m(f):=\sup_{g\in\mathcal{C}_{1}}\left(\Re\int_{\mathbb{S}^{2}}f\bar{g}\mathrm{d}\sigma(\theta)\right)^{2},\quad\mathcal{C}_{1}:=\left\{g\in\mathcal{C}:\|g\|_{2}=1\right\}.$
And for each function $f\in L^{2}(\mathbb{S}^{2})$, its constant distance $\mathrm{dist}(f,\mathcal{C})$ can be attained.
Proof of Lemma 3.1.
The first part of this lemma follows from a direct expansion, and the second part can be shown by applying an approximate identity argument [Gra14, Section 1.2.4]. For the sake of brevity, we omit the detailed proof here. Further details can be found in [K2̈5, Lemma 2.2] and [DEF+25, Lemma 3.3]. ∎
Proof of Proposition 1.12.
First, by the sphere profile decomposition as shown in [CS12, FLS16, FS24], one can see that
(21) $\|f_{y}\|_{L^{2}(\mathbb{S}^{2})}^{2}=2\|1\|_{L^{2}(\mathbb{S}^{2})}^{2}+o_{|y|\to\infty}(1)=8\pi+o_{|y|\to\infty}(1),$
and
(22) $\|\widehat{f_{y}\sigma}\|_{L^{4}(\mathbb{R}^{3})}^{4}=2\mathbf{M}^{4}\|1\|_{L^{2}(\mathbb{S}^{2})}^{4}+o_{|y|\to\infty}(1)=512\pi^{6}+o_{|y|\to\infty}(1).$
For the item $\mathrm{dist}(f,\mathcal{C})^{2}$, by the constant distance representation Lemma 3.1, we investigate
$m(f_{y})=\sup_{x\in\mathbb{R}^{3}}\left(\int_{\mathbb{S}^{2}}f_{y}(\theta)e^{ix\theta}\mathrm{d}\sigma(\theta)\right)^{2}\Big/\|e^{ix\theta}\|_{L^{2}(\mathbb{S}^{2})}^{2}=\frac{1}{4\pi}\sup_{x\in\mathbb{R}^{3}}\left(\int_{\mathbb{S}^{2}}\left(1+e^{iy\theta}\right)e^{ix\theta}\mathrm{d}\sigma(\theta)\right)^{2}.$
By a direct computation, we conclude
$m(f_{y})=4\pi\sup_{x\in\mathbb{R}^{3}}\left(\frac{\sin|x|}{|x|}+\frac{\sin|x+y|}{|x+y|}\right)^{2}.$
Then by an elementary analysis, one can obtain
$\lim_{|y|\to\infty}m(f_{y})=4\pi.$
Hence, due to Lemma 3.1, we obtain the limit behavior estimate
(23) $\mathrm{dist}(f_{y},\mathcal{C})^{2}=4\pi+o_{|y|\to\infty}(1).$
Combining the estimates (21) and (22) and (23), one can see that
$\frac{\mathbf{M}^{2}\left\|f_{y}\right\|_{L^{2}(\mathbb{S}^{2})}^{2}-\left\|\widehat{f_{y}\sigma}\right\|_{L^{4}(\mathbb{R}^{3})}^{2}}{\mathrm{dist}(f_{y},\mathcal{C})^{2}}=(2-\sqrt{2})4\pi^{2}+o_{|y|\to\infty}(1).$
This gives the desired result and completes the proof. ∎
3.3. Compactness of minimizing sequences
In this subsection, based on the spectral-gap result Proposition 1.11 and two-peak vanishing result Proposition 1.12, we show the desired existence of minimizers for the two-dimensional sphere stability constant $C_{**}$.
Proof of Theorem 1.1.
The strategy is similar to the [K2̈5, Section 4, Proof of Theorem 1.2] and the Proof of Theorem 1.3 in the previous Section 2.3. For simplicity, we only show the main ideas and omit the details. In this two-dimensional sphere case, we can obtain the unconditional result since we can establish the strict inequality in Proposition 1.11. Therefore, by following the outline in the Proof of Theorem 1.3, it remains only to check that
$\mathbf{M}^{2}-C_{SG*}=4\pi^{2}-\frac{8\pi^{2}}{5}=\frac{12\pi^{2}}{5}>0,$
which essentially can give the nonzero weak limit of minimizing sequences
$f_{n}\rightharpoonup f_{0}\;\;\text{in}\;\;L^{2}(\mathbb{S}^{2}),\quad f_{0}\neq 0.$
Then the desired existence of minimizers follows from standard arguments. ∎